Parker is 50 and wants to retire in 15 years. His family has a history of living well into their 90s. Therefore, he estimates that he will live to age 95. He currently has a salary of $120,000 and expects that he will need about 65% of that amount annually if he were retired. He can earn 9 percent in his portfolio while he is working. However, he expects that he will only earn 7 percent in his portfolio during retirement. He expects inflation to continue at 3 percent. Parker currently has $350,000 invested for his retirement. His Social Security benefit in today’s dollars is $30,000 per year at normal age retirement of age 67. His Social Security benefit will be reduced by 6 2/3 percent for each year he begins collecting before full age retirement. How much does he need to save each year to meet his retirement goals?
Answer #1
Annual amt. required after 15 years,ie. From end of 16 th year(age 65) to end of age 95th year-- for a period of 30 years = 120000*65%= $ 78000
So, total amt. to be accumulated at end of 65 years(15 years from now) ,
at a rate of 7% with an inflation rate (growing annuity)of 3%
Using PV of gowing annuity formula,
we find the PV at end of year 65 , with the above inputs:
PV(GA)=P/(r-g)*(1-((1+g)/(1+r))^n)
1…...ie.PV(GA)=78000/(7%-3%)*(1-((1+3%)/(1+7%))^30)
1328218
PV of funds available with him ,now
2..Retiremet investment = 350000
PV of social security benefits of $ 30000 p.a.
for a period of 17 years(67-50)
at an opportunity cost- rate of 9% at a decreasing annuity of
6&2/3%-3%= 3.67% net of inflation
Using the same (above) formula,
we get the PV of decreasing social benefits as
3.....PV(Decr.A)=30000/(9%-3.67%)*(1-((1-3.67%)/(1+9%))^17)
493972
4…..Total PV of funds available with him now,
350000+493972=
843972
5..Total PV of amt. to be saved to meet retirement goals=
Amts. got above as per 1-4
ie. 1328218-843972=
484246
So, this $ 484246 is the FV of growing annuity
at end of 15 years
at an opportunity cost- rate of 9% & inflation growth rate of 3%
Using the formual
FV(GA)=P*(((1+r)^n-(1+g)^n)/(r-g))
ie.484246=P*(((1+9%)^15-(1+3%)^15)/(9%-3%))
solving for P, the annual savings, we get,
$13,938
So the ANSWER is:
he need to save $ 13938 each year to meet his retirement goals